# Consistency of Neural Causal Partial Identification

### v1

@@ D.3 Counterexample
We compare the confidence intervals of the unregularized neural casual method and the Lipschitz regularized one. We choose the layerwise Lispchitz constants upper bound to be $5$. As a benchmark, we also include the bound obtained by the Double Machine Learning estimator using the EconML package. The result is shown in Figure 8. For this identifiable case example, the double ML estimator produces better bounds for the ATE. The intervals given by regularized and unregularized NCM are similar with the regularized one slightly better in this example. We conjecture that part of the reason lies in the optimization process. The landscape of the feasible region of the optimization problems (6) may be several disconnected regions. Once the algorithm chooses an initiate point in one bounded region, it will not ’jump’ from that region to other regions and stay in this region, implicitly imposing a regularization on the Lipschitz constants.
Figure 8: Comparison of Lipschitz regularized and unregularized neural causal algorithm. We use the projected gradient to regularize the weight of the neural network. For each sample size, we repeat the experiment 5 times and take the average of the upper (lower) bound.
@@ Appendix E Technical Lemmas
Lemma 3.
Let $\mu,\hat{\mu}$ be two measures on compact set $K\subset\mathbb{R}^{d_{1}}$, for any measurable functions $F,\hat{F}:K\rightarrow\mathbb{R}^{d_{2}}$, if $F$ is $L$-Lipschitz continuous, then
$W(F_{\#}\mu,\hat{F}_{\#}\hat{\mu})\leqslant LW(\mu,\hat{\mu})+\|F_{1}-F_{2}\|_{\infty}.$
@@
$\mathbb{E}_{X\sim F_{\#}\mu}[g(X)]-\mathbb{E}_{X\sim\hat{F}_{\#}\hat{\mu}}[g(X)]$ $=\mathbb{E}_{X\sim\mu}[g\circ F(X)]-\mathbb{E}_{X\sim\hat{\mu}}[g\circ\hat{F}(X)]$
$=(\mathbb{E}_{X\sim\mu}[g\circ F(X)]-\mathbb{E}_{X\sim\hat{\mu}}[g\circ F(X)])$
$\quad+(\mathbb{E}_{X\sim\hat{\mu}}[g\circ F(X)]-\mathbb{E}_{X\sim\hat{\mu}}[g\circ\hat{F}(X)]).$
@@
$\mathbb{E}_{X\sim\mu}[g\circ F(X)]-\mathbb{E}_{X\sim\hat{\mu}}[g\circ F(X)]$ $\leqslant LW(\mu,\hat{\mu}).$ (32)
@@
$\mathbb{E}_{X\sim\hat{\mu}}[g\circ F(X)]-\mathbb{E}_{X\sim\hat{\mu}}[g\circ\hat{F}(X)]$ $\leqslant\mathbb{E}_{X\sim\hat{\mu}}[|g\circ F(X)-g\circ\hat{F}(X)|]$
$\leqslant\mathbb{E}_{X\sim\hat{\mu}}[\|F(X)-\hat{F}(X)\|]$
$\leqslant\|F(X)-\hat{F}(X)\|_{\infty}.$ (33)
Combine (32) and (33), we have the conclusion. ∎
Lemma 4.
Let $\mu_{1},\cdots,\mu_{n},\hat{\mu}_{1},\cdots,\hat{\mu}_{n}$ be measure on $\mathbb{R}^{d}$, for any $p_{1},\cdots,p_{n}\in[0,1]$ such that $\sum_{k=1}^{n}p_{k}=1$, we have
$W\left(\sum_{k=1}^{n}p_{k}\mu_{k},\sum_{k=1}^{n}p_{k}\hat{\mu}_{k}\right)\leqslant\sum_{k=1}^{n}p_{k}W(\mu_{k},\hat{\mu}_{k}).$
@@
$\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\mu_{k}}[f(X)]-\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\hat{\mu}_{k}}[f(X)]$ $=\sum_{k=1}^{n}p_{k}(\mathbb{E}_{X_{k}\sim\mu_{k}}[f(X_{k})]-\mathbb{E}_{X_{k}\sim\hat{\mu}_{k}}[f(X_{k})])$
$\leqslant\sum_{k=1}^{n}p_{k}W(\mu_{k},\hat{\mu}_{k}).$
@@
$W\left(\sum_{k=1}^{n}p_{k}\mu_{k},\sum_{k=1}^{n}p_{k}\mu_{k}\right)$ $=\sup_{f\ \text{1-Lipschitz}}\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\mu_{k}}[f(X)]-\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\hat{\mu}_{k}}[f(X)]$
$\leqslant\sum_{k=1}^{n}p_{k}W(\mu_{k},\hat{\mu}_{k}).$
Lemma 5.
Given two measures $\mu=\mu_{1}\otimes\cdots\otimes\mu_{n},\hat{\mu}=\hat{\mu}_{1}\otimes\cdots\otimes\hat{\mu}_{n}$, we have
$W(\mu,\hat{\mu})\leqslant\sum_{k=1}^{n}W(\mu_{k},\hat{\mu}_{k}).$
@@
For any $\epsilon>0$, let $\pi_{k}$ be a coupling of $\mu_{k},\hat{\mu}_{k}$ such that
$\mathbb{E}_{X_{k},Y_{k}\sim\pi_{k}}[\|X_{k}-Y_{k}\|_{1}]\leqslant W(\mu_{k},\hat{\mu}_{k})+\epsilon.$
@@
$W(\mu,\hat{\mu})$ $\leqslant\mathbb{E}_{\pi_{1}\otimes\cdots\otimes\pi_{n}}[\|X-Y\|]$
@@
Lemma 6 (Approximation error of Gumbel-Softmax layer).
@@
Without loss of generality, we may assume that $\sum_{k=1}^{n}p_{k}=1$. Otherwise, we can divide the denominator and numerator in (34) by $\sum_{k=1}^{n}p_{k}$. Let $\Delta_{n}=\left\{(x_{1},\cdots,x_{n}):\sum_{k=1}^{n}x_{k}=1\right\}$. We construct a transportation map $T:\Delta_{n}\rightarrow\Delta_{n}$ as follows. $T(x_{1},\cdots,x_{n})$ is a $n$-dimensional vector with all zeros except that the $i=\arg\max_{j}x_{j}$-th entry being one. If there are multiple coordinates that is maximum, we choose the first coordinate and let the remaining coordinate be zero. It is easy to verify that $T_{\#}\mu^{\tau}$ and $\mu^{0}$ have the same distribution [35]. For any $\delta>0$, we have
@@
Lemma 7 (Approximation Error of Sinkhorn distance).
@@
Let $h(T)=-\sum_{i,j=1}^{n,m}T_{ij}\log T_{ij}+1$, by [34, Proposition 1], we have
$0\leqslant W_{1,\lambda}(\mu,\nu)-W(\mu,\nu)$ $\leqslant c\lambda,$ (35)

### v2

@@ D.3 Counterexample
We compare the confidence intervals of the unregularized neural casual method and the Lipschitz regularized one under different architectures. We choose the layerwise Lispchitz constants upper bound to be $5$ and $1.2$. As a benchmark, we also include the bound obtained by the Double Machine Learning estimator using the EconML package. The result is shown in Figure 12. For this identifiable case example, the double ML estimator produces better bounds for the ATE. The intervals given by regularized and unregularized NCM are similar to the regularized one slightly better in the left figure, where we use medium-sized NNs. However, in the right figure, the obtained intervals after regularization are tighter, although slightly biased, compared with the regularized setting. From these two experiments, we conclude that the architecture of NNs will also influence the results and adding regularization during the training process can prevent extreme confidence intervals or inconsistency.
Figure 12: Comparison of Lipschitz regularized and unregularized neural causal algorithm. The two figures show the results of different architectures. The figure on the left side uses a medium-sized NN (width 128, depth 3) to approximate each structural function, while the right figure uses extremely small NNs (width 3, depth 1). In all experiments, we use the projected gradient to regularize the weight of the neural network. For each sample size, we repeat the experiment 5 times and take the average of the upper (lower) bound.
@@ Appendix E Technical Lemmas
Lemma E.1.
Let $\mu,\widehat{\mu}$ be two measures on compact set $K\subset\mathbb{R}^{d_{1}}$, for any measurable functions $F,\widehat{F}:K\rightarrow\mathbb{R}^{d_{2}}$, if $F$ is $L$-Lipschitz continuous, then
$W(F_{\#}\mu,\widehat{F}_{\#}\widehat{\mu})\leqslant LW(\mu,\widehat{\mu})+\|F_{1}-F_{2}\|_{\infty}.$
@@
$\mathbb{E}_{X\sim F_{\#}\mu}[g(X)]-\mathbb{E}_{X\sim\widehat{F}_{\#}\widehat{\mu}}[g(X)]$ $=\mathbb{E}_{X\sim\mu}[g\circ F(X)]-\mathbb{E}_{X\sim\widehat{\mu}}[g\circ\widehat{F}(X)]$
$=(\mathbb{E}_{X\sim\mu}[g\circ F(X)]-\mathbb{E}_{X\sim\widehat{\mu}}[g\circ F(X)])$
$\quad+(\mathbb{E}_{X\sim\widehat{\mu}}[g\circ F(X)]-\mathbb{E}_{X\sim\widehat{\mu}}[g\circ\widehat{F}(X)]).$
@@
$\mathbb{E}_{X\sim\mu}[g\circ F(X)]-\mathbb{E}_{X\sim\widehat{\mu}}[g\circ F(X)]$ $\leqslant LW(\mu,\widehat{\mu}).$ (E.1)
@@
$\mathbb{E}_{X\sim\widehat{\mu}}[g\circ F(X)]-\mathbb{E}_{X\sim\widehat{\mu}}[g\circ\widehat{F}(X)]$ $\leqslant\mathbb{E}_{X\sim\widehat{\mu}}[|g\circ F(X)-g\circ\widehat{F}(X)|]$
$\leqslant\mathbb{E}_{X\sim\widehat{\mu}}[\|F(X)-\widehat{F}(X)\|]$
$\leqslant\|F(X)-\widehat{F}(X)\|_{\infty}.$ (E.2)
Combine (E.1) and (E.2), we have the conclusion. ∎
Lemma E.2.
Let $\mu_{1},\cdots,\mu_{n},\widehat{\mu}_{1},\cdots,\widehat{\mu}_{n}$ be measure on $\mathbb{R}^{d}$, for any $p_{1},\cdots,p_{n}\in[0,1]$ such that $\sum_{k=1}^{n}p_{k}=1$, we have
$W\left(\sum_{k=1}^{n}p_{k}\mu_{k},\sum_{k=1}^{n}p_{k}\widehat{\mu}_{k}\right)\leqslant\sum_{k=1}^{n}p_{k}W(\mu_{k},\widehat{\mu}_{k}).$
@@
$\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\mu_{k}}[f(X)]-\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\widehat{\mu}_{k}}[f(X)]$ $=\sum_{k=1}^{n}p_{k}(\mathbb{E}_{X_{k}\sim\mu_{k}}[f(X_{k})]-\mathbb{E}_{X_{k}\sim\widehat{\mu}_{k}}[f(X_{k})])$
$\leqslant\sum_{k=1}^{n}p_{k}W(\mu_{k},\widehat{\mu}_{k}).$
@@
$W\left(\sum_{k=1}^{n}p_{k}\mu_{k},\sum_{k=1}^{n}p_{k}\mu_{k}\right)$ $=\sup_{f\ \text{1-Lipschitz}}\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\mu_{k}}[f(X)]-\mathbb{E}_{X\sim\sum_{k=1}^{n}p_{k}\widehat{\mu}_{k}}[f(X)]$
$\leqslant\sum_{k=1}^{n}p_{k}W(\mu_{k},\widehat{\mu}_{k}).$
Lemma E.3.
Given two measures $\mu=\mu_{1}\otimes\cdots\otimes\mu_{n},\widehat{\mu}=\widehat{\mu}_{1}\otimes\cdots\otimes\widehat{\mu}_{n}$, we have
$W(\mu,\widehat{\mu})\leqslant\sum_{k=1}^{n}W(\mu_{k},\widehat{\mu}_{k}).$
@@
For any $\epsilon>0$, let $\pi_{k}$ be a coupling of $\mu_{k},\widehat{\mu}_{k}$ such that
$\mathbb{E}_{X_{k},Y_{k}\sim\pi_{k}}[\|X_{k}-Y_{k}\|_{1}]\leqslant W(\mu_{k},\widehat{\mu}_{k})+\epsilon.$
@@
$W(\mu,\widehat{\mu})$ $\leqslant\mathbb{E}_{\pi_{1}\otimes\cdots\otimes\pi_{n}}[\|X-Y\|]$
@@
Lemma E.4 (Approximation error of Gumbel-Softmax layer).
@@
Without loss of generality, we may assume that $\sum_{k=1}^{n}p_{k}=1$. Otherwise, we can divide the denominator and numerator in (E.3) by $\sum_{k=1}^{n}p_{k}$. Let $\Delta_{n}=\left\{(x_{1},\cdots,x_{n}):\sum_{k=1}^{n}x_{k}=1\right\}$. We construct a transportation map $T:\Delta_{n}\rightarrow\Delta_{n}$ as follows. $T(x_{1},\cdots,x_{n})$ is a $n$-dimensional vector with all zeros except that the $i=\arg\max_{j}x_{j}$-th entry being one. If there are multiple coordinates that is maximum, we choose the first coordinate and let the remaining coordinate be zero. It is easy to verify that $T_{\#}\mu^{\tau}$ and $\mu^{0}$ have the same distribution [36]. For any $\delta>0$, we have
@@
Lemma E.5 (Approximation Error of Sinkhorn distance).
@@
Let $h(T)=-\sum_{i,j=1}^{n,m}T_{ij}\log T_{ij}+1$, by [35, Proposition 1], we have
$0\leqslant W_{1,\lambda}(\mu,\nu)-W(\mu,\nu)$ $\leqslant c\lambda,$ (E.4)